The difference in hours between​ full-timers and the​ part-timers who work 6 hours a day is 5 working hours. How many hours per day do​ full-timers work?


1
Expert's answer
2021-04-15T07:32:51-0400

Solution.

If the difference is 5 hours, full-timers work

6 + 5 = 11 hours.

Considering most people work for 5 days, part-timers work 5 * 6= 30 hours for week. Then full-timers work 5 * 11=55 hours for week.

Answer. 11 hours.